OpenStudy (bohotness):

What a bag containing black and white marbles, of the marbles are black. Which could be the numbers of black and white marbles in the bag? A. 3 black and 8 white B. 12 black and 20 white C. 20 black and 12 white D. 8 black and 3 white

OpenStudy (anonymous):

missing something?

OpenStudy (bohotness):

In a bag containing black and white marbles, of the marbles are black. Which could be the numbers of black and white marbles in the bag? A. 3 black and 8 white B. 12 black and 20 white C. 20 black and 12 white D. 8 black and 3 white

OpenStudy (bohotness):

3/8

OpenStudy (anonymous):

the ratio of black marbles to the total must be \(3:8\) or \(\frac{3}{8}\)

OpenStudy (bohotness):

Oh

OpenStudy (anonymous):

not A because \[\frac{3}{3+8}=\frac{3}{11}\neq \frac{3}{8}\]

OpenStudy (anonymous):

maybe B try \[\frac{12}{12+20}\]

OpenStudy (bohotness):

Ok

OpenStudy (bohotness):

But b is 0.375

OpenStudy (bohotness):

@satellite73